# Priors from Envisioned Posterior Judgments: A Novel Elicitation Approach With Application to Bayesian Clinical Trials

### v2

@@ 5 Discussion
We have described and pilot-tested a new prior elicitation approach that addresses two hurdles to obtaining credible prior distributions. First, we showed how our process for inferring the prior distribution from eliciting judgments about an expert’s envisioned posterior quantities given hypothetical outcome data also provides an assessment of prior-posterior coherence, which can then be used by the experts to improve the coherency of their judgments. (We emphasize, however, that while prior-posterior coherence is necessary for a prior distribution to be credible, it should not be viewed as a sufficient condition.) Second, through asking experts for judgments about expected values (i.e., point estimates), rather than probabilistic statements about uncertainty, we reduced their cognitive burden and avoided potential errors that might arise from misunderstanding of probability concepts. The case study results provide a proof-in-principle demonstration of the feasibility of the proposed approach.
@@
Although our results are promising, considerable additional development will be needed to make this approach broadly and easily accessible. First, the technical framework needs to be extended. In our case study, we considered only the elicitation of a single mean parameter from a Normal data model and assumed that the expert’s prior uncertainty could be captured using a conjugate Normal distribution. This provided an approximate analytical expression for the posterior distribution and simplified the computations. However, for more complex problems involving multiple parameters or other types of outcomes (e.g., survival outcomes), conjugate priors will be unavailable and will require the use of more complex computational algorithms (based on Markov Chain Monte Carlo techniques, for example) to obtain posterior quantities [26]. In addition, further exploration is needed into whether RMSD is the most appropriate metric for representing the degree of coherency.
@@
Second, elicitation is not purely a mathematical exercise but is a process in which the results are strongly influenced by how the process is implemented. At this time, the optimal elicitation process remains unclear on multiple facets. Small changes to the presentation of materials can impact the outcomes. As an example, during the revision stage, we changed the presentation of our scenarios from stand-alone questions only and added a table displaying the responses from all scenarios. Several experts commented that this helped them to see much more easily where their responses were not coherent, and assisted them when revising their judgments. Other observations raise similar questions. For example, how much guidance should be given to the experts before the initial elicitation? The poor coherence obtained in the initial elicitation suggests that the experts could have benefited from more guidance than was given in this study. Yet, giving too much information initially could overwhelm the experts leading them to be uncomfortable with making any judgments. Nonetheless, the marked improvement achieved in the revised elicitation suggests that providing experts with a summary of common issues (like those identified in this study) might help them achieve a higher level of coherence from the outset. Obtaining reasonable responses at the outset is important to pre-empting anchoring bias, wherein experts are reluctant to revise initial responses even when they are shown to be incoherent. Some experts reported feeling that they were being “pressured” to revise their initial responses rather than recognizing it as an opportunity to reflect on whether their initial responses had captured their beliefs accurately. More consultation with experts on how feedback should be presented and on the choice of language is needed to determine how to avoid such cognitive biases. Addressing these questions is pivotal for refining our methodology and improving the overall quality of the research outcomes. Future research may want to focus on these aspects and provide more structured elicitation methods for more robust and reliable results. Last, one challenge of elicitation is a lack of practical tools that can handle the mathematical details of model fitting [27, 11].
@@
In this study, we did not seek to compare the results of our new approach to what prior distributions would have been obtained if an existing elicitation approach had been used. While such comparisons would provide some indication of the sensitivity of elicited priors to the elicitation approach, judging whether one approach is “better” than another is inherently difficult because no standard criteria for making such comparisons have been established. For example, one might consider assessing prior-posterior coherence for a prior distribution obtained by direct elicitation by eliciting posterior judgments as described here and then computing the coherency metric associated with this directly elicited prior. But of course, this will always show poorer coherence than our approach since our procedure by construction selects the prior that maximizes coherency.

### v3

@@ 4 Discussion
We have described and pilot-tested a new prior elicitation approach that addresses two hurdles to obtaining credible prior distributions. First, we showed how our process for inferring the prior distribution from eliciting judgments about an expert’s envisioned posterior quantities given hypothetical outcome data also provides an assessment of prior-posterior coherence, which can then be used by the experts to improve the coherency of their judgments. Second, through asking experts for judgments about expected values (i.e., point estimates), rather than probabilistic statements about uncertainty, we reduced their cognitive burden and avoided potential errors that might arise from misunderstanding of probability concepts. The case study results provide a proof-in-principle demonstration of the feasibility of the proposed approach.
@@
Although our results are promising, considerable additional development will be needed to make this approach broadly and easily accessible. First, the technical framework needs to be extended. In our case study, we considered only the elicitation of a single mean parameter from a normal data model and assumed that the expert’s prior uncertainty could be captured using a conjugate normal distribution. This provided an approximate analytical expression for the posterior distribution and simplified the computations. However, for more complex problems involving multiple parameters or other types of outcomes (e.g., survival outcomes), conjugate priors will be unavailable and will require the use of more complex computational algorithms (e.g., Markov Chain Monte Carlo techniques, for example) to obtain posterior quantities [27]. In addition, further exploration is needed into whether RMSD is the most appropriate metric for representing the degree of coherency and what constitutes adequate coherency.
@@
Second, elicitation is not purely a mathematical exercise but is a process in which the results are strongly influenced by how the process is implemented. At this time, the optimal elicitation process remains unclear on multiple facets. Small changes to the presentation of materials can impact the outcomes. As an example, during the revision stage, we changed the presentation of our scenarios from stand-alone questions only and added a table displaying the responses from all scenarios. Several experts commented that this helped them to see much more easily where their responses were not coherent, and assisted them when revising their judgments. Other observations raise similar concerns. For example, how much guidance should be given to the experts before the initial elicitation? The poor coherence obtained in the initial elicitation suggests that the experts could have benefited from more guidance than was given in this study. Yet, giving too much information initially could overwhelm the experts leading them to be uncomfortable with making any judgments. Nonetheless, the marked improvement achieved in the revised elicitation suggests that providing experts with a summary of common issues (like those identified in this study) might help them achieve a higher level of coherence from the outset. Obtaining reasonable responses at the outset is important to pre-empting anchoring bias, wherein experts are reluctant to revise initial responses even when they are shown to be incoherent. Messaging to experts that their initial judgments should be treated as ”tentative” might have pre-empted reports from some experts of feeling pressured to revise their initial responses, and clearer communication that the revision process is meant to provide an opportunity to reflect on whether their initial responses had captured their beliefs accurately might have made the task less stressful to the experts. More consultation with experts on how feedback should be presented and on the choice of language is needed to determine how to avoid such cognitive biases. Addressing these questions is pivotal for refining our methodology and improving the overall quality of the research outcomes. Future research may want to focus on these aspects and provide more structured elicitation methods for more robust and reliable results. Last, one challenge of elicitation is a lack of practical tools for fitting the appropriate mathematical models [28, 11].
@@
In this study, we did not seek to compare the results of our new approach with what prior distributions would have been obtained if another existing elicitation approach had been used. While such comparisons would provide some indication of the sensitivity of elicited priors to the elicitation approach, judging whether one approach is “better” than another is inherently difficult because no standard criteria for making such comparisons have been established. For example, one might consider assessing prior-posterior coherence for a prior distribution obtained by direct elicitation by eliciting posterior judgments as described here and then computing the coherency metric associated with this directly elicited prior. But, of course, this will always show poorer coherence than our approach since our procedure by construction selects the prior that maximizes coherency.
While we have illustrated our approach within the context of a stepped-wedge cluster randomized trial, the proposed method is applicable to any context that requires elicitation of a prior distribution, such as for other clinical trial designs or even in non-trial settings, such as health economics, where informed decision-making under uncertainty is essential. In particular, expert elicitation can be especially useful in early-phase trials, where direct evidence is limited and sample sizes are small [29, 30, 31]. For example, many targeted and immunotherapies are now applied across multiple cancer types due to shared molecular alterations. In these settings, Bayesian elicitation is a particularly valuable tool, allowing the formal integration of expert opinion and historical evidence to guide dose selection, endpoint estimation, and trial design, ultimately improving efficiency and reducing uncertainty.
5 Conclusions
@@
List of abbreviations
PfP: Prior from posterior
RMSD: Root mean square discrepancy
WnW: Walk ’n Watch
SW-CRT: Stepped-wedge cluster randomized trials
MWT: Minute walk test
Declarations
Ethics approval and consent to participate
Not Applicable
Consent for publication
All authors have reviewed and aggreed on the final version of the manuscript
Availability of data and materials
No new data was generated for this study
Competing interests
All authors declare no competing interests
Funding
The CanStroke Recovery Trials Platform is supported by the Canada Brain Research Fund (CBRF), an innovative arrangement between the Government of Canada (through Health Canada) and Brain Canada Foundation and the Heart and Stroke Foundation Canadian Partnership for Stroke Recovery, University of Calgary, and the Sunnybrook Research Institute.
Authors’ contributions
HW and YO both conceived and developed the idea. HW contributed to the manuscript’s revision, while YO led the manuscript writing. JE is the PI of the case study and facilitated the elicitation between statisticians (HW and YO) and experts (The WnW Research Team). JE and DZ provided significant contributions to the manuscript development. The WnW Research Team participated in the elicitation exercise.
Acknowledgments
The authors gratefully acknowledge the support and collaboration of the CanStroke Recovery Trials Platform.
Supplementary Materials
S1: Walk n Watch secular trend elicitation questionnaire
Name:
The purpose of this exercise is to elicit your judgments about how much the efficacy of usual care changed during the Walk ’n Watch trial from year 2021 to year 2022. Many things could have changed the efficacy, including improvements in practice (positive) or staffing shortages (negative). You will be asked to state your estimate of the true change in efficacy in the study population (patients eligible for the Walk ‘n Watch trial) given various hypothetical outcomes in a sample of data. That is, we are interested in what you would conclude if the hypothetical outcomes presented here had been the actual data observed in the trial. (For this exercise, please disregard any actual trial data that you may have seen.)
There is no ”right answer” – rather, your response should reflect an assessment that combines both (a) your prior beliefs derived from past clinical data and experiences over 2021 and 2022, and (b) the sample data, according to how much weight you give to each component. For example, in one of the scenarios below, the sample data show an increase of 30 metres based on 10 patients in 2021 and 10 patients in 2022. But if you felt strongly a priori that the true change is near zero, say, and that this magnitude of change seen in the data is implausibly large (and perhaps primarily reflecting noise arising from the small sample), you should feel free to give little weight to the sample data and to choose a much smaller change or no change (zero) as your response. In contrast, in the scenario where the same 30 metres increase is based on 100 patients in 2021 and 100 patients in 2022, you may wish to give more weight to the data, and your response should be adjusted accordingly. We will ask you to consider 16 scenarios involving varying amounts of data and a variety of outcomes.
Scenarios
For the first scenario, assume you have no sample data.
  1. 1.
    By how much do you think the average improvement changed in the study population (that is, stroke patients eligible for the Walk ‘n Watch study) from 2021 to 2022?
    Response:                  metres
    Please note, 0 metres would indicate that the clinical practice and improvements in patients was similar from 2021 to 2022; a positive value would indicate that practice had improved in 2022 resulting in an increase of the amount of metres in the 6MWT; a negative value would indicate that the practice had deteriorated in 2022 resulting in a reduction in the 6MWT.
For the next 5 scenarios, the sample consists of 10 stroke patients from each year.
  1. 2.
    Suppose that in your random sample of 10 usual care stroke patients in 2021, their average improvement in 6MWT was 140 metres. Suppose also that in your random sample of 10 usual care stroke patients in 2022, their average improvement was 140 metres (no change compared to 2021).
    By how much do you think the average improvement changed in the study population (that is, stroke patients eligible for the Walk ‘n Watch study) from 2021 to 2022?
    Response:                  metres
    Please note, 0 metres would indicate that the clinical practice and improvements in patients was similar from 2021 to 2022; a positive value would indicate that practice had improved in 2022 resulting in an increase of the amount of metres in the 6MWT; a negative value would indicate that the practice had deteriorated in 2022 resulting in a reduction in the 6MWT.
  2. 3.
    Suppose that in your random sample of 10 usual care stroke patients in 2021, their average improvement in 6MWT was 140 metres. Suppose also that in your random sample of 10 usual care stroke patients in 2022, their average improvement was 150 metres (10 metre greater improvement compared to 2021).
    By how much do you think the average improvement changed in your population (that is, stroke patients eligible for the Walk ‘n Watch study) from 2021 to 2022?
    Response:                  metres
    Please note, 0 metres would indicate that the clinical practice and improvements in patients was similar from 2021 to 2022; a positive value would indicate that practice had improved in 2022 resulting in an increase of the amount of metres in the 6MWT; a negative value would indicate that the practice had deteriorated in 2022 resulting in a reduction in the 6MWT.
  3. 4.
    Suppose that in your random sample of 10 usual care stroke patients in 2021, their average improvement in 6MWT was 140 metres. Suppose also that in your random sample of 10 usual care stroke patients in 2022, their average improvement was 170 metres (30 metre greater improvement compared to 2021).
    By how much do you think the average improvement changed in your population (that is, stroke patients eligible for the Walk ‘n Watch study) from 2021 to 2022?
    Response:                  metres
    Please note, 0 metres would indicate that the clinical practice and improvements in patients was similar from 2021 to 2022; a positive value would indicate that practice had improved in 2022 resulting in an increase of the amount of metres in the 6MWT; a negative value would indicate that the practice had deteriorated in 2022 resulting in a reduction in the 6MWT.
  4. 5.
    Suppose that in your random sample of 10 usual care stroke patients in 2021, their average improvement in 6MWT was 140 metres. Suppose also that in your random sample of 10 usual care stroke patients in 2022, their average improvement was 130 metres (10 metre reduction compared to 2021).
    By how much do you think the average improvement changed in your population (that is, stroke patients eligible for the Walk ‘n Watch study) from 2021 to 2022?
    Response:                  metres
    Please note, 0 metres would indicate that the clinical practice and improvements in patients was similar from 2021 to 2022; a positive value would indicate that practice had improved in 2022 resulting in an increase of the amount of metres in the 6MWT; a negative value would indicate that the practice had deteriorated in 2022 resulting in a reduction in the 6MWT.
  5. 6.
    Suppose that in your random sample of 10 usual care stroke patients in 2021, their average improvement in 6MWT was 140 metres. Suppose also that in your random sample of 10 usual care stroke patients in 2022, their average improvement was 110 metres (30 metre reduction compared to 2021).
    By how much do you think the average improvement changed in your population (that is, stroke patients eligible for the Walk ‘n Watch study) from 2021 to 2022?
    Response:                  metres
    Please note, 0 metres would indicate that the clinical practice and improvements in patients was similar from 2021 to 2022; a positive value would indicate that practice had improved in 2022 resulting in an increase of the amount of metres in the 6MWT; a negative value would indicate that the practice had deteriorated in 2022 resulting in a reduction in the 6MWT.
For the next 5 scenarios, the sample consists of 30 stroke patients from each year.
  1. 7.
    Suppose that in your random sample of 30 usual care stroke patients in 2021, their average improvement in 6MWT was 140 metres. Suppose also that in your random sample of 30 usual care stroke patients in 2022, their average improvement was 140 metres (no change compared to 2021).
    By how much do you think the average improvement changed in the study population (that is, stroke patients eligible for the Walk ‘n Watch study) from 2021 to 2022?
    Response:                  metres
    Please note, 0 metres would indicate that the clinical practice and improvements in patients was similar from 2021 to 2022; a positive value would indicate that practice had improved in 2022 resulting in an increase of the amount of metres in the 6MWT; a negative value would indicate that the practice had deteriorated in 2022 resulting in a reduction in the 6MWT.
  2. 8.
    Suppose that in your random sample of 30 usual care stroke patients in 2021, their average improvement in 6MWT was 140 metres. Suppose also that in your random sample of 30 usual care stroke patients in 2022, their average improvement was 150 metres (10 metre greater improvement compared to 2021).
    By how much do you think the average improvement changed in your population (that is, stroke patients eligible for the Walk ‘n Watch study) from 2021 to 2022?
    Response:                  metres
    Please note, 0 metres would indicate that the clinical practice and improvements in patients was similar from 2021 to 2022; a positive value would indicate that practice had improved in 2022 resulting in an increase of the amount of metres in the 6MWT; a negative value would indicate that the practice had deteriorated in 2022 resulting in a reduction in the 6MWT.
  3. 9.
    Suppose that in your random sample of 30 usual care stroke patients in 2021, their average improvement in 6MWT was 140 metres. Suppose also that in your random sample of 30 usual care stroke patients in 2022, their average improvement was 170 metres (30 metre greater improvement compared to 2021).
    By how much do you think the average improvement changed in your population (that is, stroke patients eligible for the Walk ‘n Watch study) from 2021 to 2022?
    Response:                  metres
    Please note, 0 metres would indicate that the clinical practice and improvements in patients was similar from 2021 to 2022; a positive value would indicate that practice had improved in 2022 resulting in an increase of the amount of metres in the 6MWT; a negative value would indicate that the practice had deteriorated in 2022 resulting in a reduction in the 6MWT.
  4. 10.
    Suppose that in your random sample of 30 usual care stroke patients in 2021, their average improvement in 6MWT was 140 metres. Suppose also that in your random sample of 30 usual care stroke patients in 2022, their average improvement was 130 metres (10 metre reduction compared to 2021).
    By how much do you think the average improvement changed in your population (that is, stroke patients eligible for the Walk ‘n Watch study) from 2021 to 2022?
    Response:                  metres
    Please note, 0 metres would indicate that the clinical practice and improvements in patients was similar from 2021 to 2022; a positive value would indicate that practice had improved in 2022 resulting in an increase of the amount of metres in the 6MWT; a negative value would indicate that the practice had deteriorated in 2022 resulting in a reduction in the 6MWT.
  5. 11.
    Suppose that in your random sample of 30 usual care stroke patients in 2021, their average improvement in 6MWT was 140 metres. Suppose also that in your random sample of 30 usual care stroke patients in 2022, their average improvement was 110 metres (30 metre reduction compared to 2021).
    By how much do you think the average improvement changed in your population (that is, stroke patients eligible for the Walk ‘n Watch study) from 2021 to 2022?
    Response:                  metres
    Please note, 0 metres would indicate that the clinical practice and improvements in patients was similar from 2021 to 2022; a positive value would indicate that practice had improved in 2022 resulting in an increase of the amount of metres in the 6MWT; a negative value would indicate that the practice had deteriorated in 2022 resulting in a reduction in the 6MWT.
For the next 5 questions, the sample consists of 100 stroke patients from each year.
  1. 12.
    Suppose that in your random sample of 100 usual care stroke patients in 2021, their average improvement in 6MWT was 140 metres. Suppose also that in your random sample of 100 usual care stroke patients in 2022, their average improvement was 140 metres (no change compared to 2021).
    By how much do you think the average improvement changed in the study population (that is, stroke patients eligible for the Walk ‘n Watch study) from 2021 to 2022?
    Response:                  metres
    Please note, 0 metres would indicate that the clinical practice and improvements in patients was similar from 2021 to 2022; a positive value would indicate that practice had improved in 2022 resulting in an increase of the amount of metres in the 6MWT; a negative value would indicate that the practice had deteriorated in 2022 resulting in a reduction in the 6MWT.
  2. 13.
    Suppose that in your random sample of 100 usual care stroke patients in 2021, their average improvement in 6MWT was 140 metres. Suppose also that in your random sample of 100 usual care stroke patients in 2022, their average improvement was 150 metres (10 metre greater improvement compared to 2021).
    By how much do you think the average improvement changed in your population (that is, stroke patients eligible for the Walk ‘n Watch study) from 2021 to 2022?
    Response:                  metres
    Please note, 0 metres would indicate that the clinical practice and improvements in patients was similar from 2021 to 2022; a positive value would indicate that practice had improved in 2022 resulting in an increase of the amount of metres in the 6MWT; a negative value would indicate that the practice had deteriorated in 2022 resulting in a reduction in the 6MWT.
  3. 14.
    Suppose that in your random sample of 100 usual care stroke patients in 2021, their average improvement in 6MWT was 140 metres. Suppose also that in your random sample of 100 usual care stroke patients in 2022, their average improvement was 170 metres (30 metre greater improvement compared to 2021).
    By how much do you think the average improvement changed in your population (that is, stroke patients eligible for the Walk ‘n Watch study) from 2021 to 2022?
    Response:                  metres
    Please note, 0 metres would indicate that the clinical practice and improvements in patients was similar from 2021 to 2022; a positive value would indicate that practice had improved in 2022 resulting in an increase of the amount of metres in the 6MWT; a negative value would indicate that the practice had deteriorated in 2022 resulting in a reduction in the 6MWT.
  4. 15.
    Suppose that in your random sample of 100 usual care stroke patients in 2021, their average improvement in 6MWT was 140 metres. Suppose also that in your random sample of 100 usual care stroke patients in 2022, their average improvement was 130 metres (10 metre reduction compared to 2021).
    By how much do you think the average improvement changed in your population (that is, stroke patients eligible for the Walk ‘n Watch study) from 2021 to 2022?
    Response:                  metres
    Please note, 0 metres would indicate that the clinical practice and improvements in patients was similar from 2021 to 2022; a positive value would indicate that practice had improved in 2022 resulting in an increase of the amount of metres in the 6MWT; a negative value would indicate that the practice had deteriorated in 2022 resulting in a reduction in the 6MWT.
  5. 16.
    Suppose that in your random sample of 100 usual care stroke patients in 2021, their average improvement in 6MWT was 140 metres. Suppose also that in your random sample of 100 usual care stroke patients in 2022, their average improvement was 110 metres (30 metre reduction compared to 2021).
    By how much do you think the average improvement changed in your population (that is, stroke patients eligible for the Walk ‘n Watch study) from 2021 to 2022?
    Response:                  metres
    Please note, 0 metres would indicate that the clinical practice and improvements in patients was similar from 2021 to 2022; a positive value would indicate that practice had improved in 2022 resulting in an increase of the amount of metres in the 6MWT; a negative value would indicate that the practice had deteriorated in 2022 resulting in a reduction in the 6MWT.
S2: Example of individual feedback
Here is an example of individual feedback for expert 9
Individual feedback
The fitted model indicates that you expect the one-year change to be a decrease of 15.8 meters with a standard deviation (SD) of 16 meters. Your responses largely are consistent with each other and with the predicted response from the model, as shown in the plot, though there are some scenarios which do not fit well. If you had been perfectly consistent, then your responses and the modeled response would fall on the straight line shown. The SD of 16 reflects the estimate of your (un)certainty about the one-year change. Roughly, it means that you are  95% sure that the true one-year change is in the range (-48, 16). (Note that range will change (likely to become narrower) if you decide to revise your responses to achieve greater consistency.)
The exact values of the points on the plot are shown in the table below. The rows highlighted in red are the scenarios in which your response (column ‘Expert’) is highly inconsistent because it falls outside the range between your initial judgment prior to seeing any data (cell in green) and the observed result in the data (column ‘Observed mean change’), and it also is far from what the value predicted by the model. We ask that you reflect on and revise your responses for these scenarios and/or your initial judgment (if you feel it no longer represents your beliefs) to resolve these discrepancies as they have a major impact on the results. The rows highlighted in yellow are scenarios that are modestly inconsistent, so revising them would improve consistency but would have less impact on the results. Examples:
In Scenario 7 (‘G’), your response was -40. However, the data suggest no change, so if you keep your initial judgment of -10, a consistent response would be between 0 and -10.
In Scenario 5 (‘E’), your response was -20. The data suggest a change of -10, which exactly supports your initial judgment, so it would make sense to choose -10 as your response. However, because your response is still somewhat close to the consistent choice, revising to -10 likely won’t have a great impact on the final result.
Scenario Plot Symbol n Observed Mean Change Expert Response Model-fitted Responses Difference
1 A NA NA -10 -15.8 5.8
2 B 10 0 -10 -13.6 3.6
3 C 10 +10 -5 -12.3 7.3
4 D 10 +30 -5 -9.5 4.5
5 E 10 -10 -20 -15 -5
6 F 10 -30 -10 -17.7 7.7
7 G 30 0 -40 -10.7 -29.3
8 H 30 +10 -5 -7.5 2.5
9 I 30 +30 0 -1.1 1.1
10 J 30 -10 -15 -13.9 -1.1
11 K 30 -30 -20 -20.4 0.4
12 L 100 0 -5 -6.1 1.1
13 M 100 +10 5 0 5
14 N 100 +30 10 12.3 -2.3
15 O 100 -10 -15 -12.2 -2.8
16 P 100 -30 -35 -24.5 -10.5